Interest rate.  Upstate Bank is offering​ long-term certificates of deposit with a face value of ​$100,000 ​(future value). Bank customers can buy these CDs today for ​$60,000 and will receive the ​$100,000 in 25 years. What interest rate is the bank paying on these​ CDs?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

100,000=60000*(1+r/100)^25

(100,000/60000)^(1/25)=(1+r/100)

(1+r/100)=1.0206

r=1.0206-1

=2.06%(Approx).